OpenStudy (anonymous):

Find the next three terms of the sequence. Then write a rule for the sequence. 648, 216, 72, 24

OpenStudy (anonymous):

keep dividing by 3 24/3 = 8

OpenStudy (anonymous):

8/3 8/9 = 8/3 * 1/3

OpenStudy (cwrw238):

each term is divided by 3 to get the next as iStutts said

OpenStudy (cwrw238):

next 3 terms are 8 , 8/3 and 8/9

OpenStudy (anonymous):

so the rule is multiply each term by 3

OpenStudy (anonymous):

divide

OpenStudy (cwrw238):

no divide

OpenStudy (anonymous):

8,2.6,0.8

OpenStudy (anonymous):

are these the next 3 numbers

OpenStudy (anonymous):

in decimal form, yes.

OpenStudy (anonymous):

should they be in decimal form

OpenStudy (cwrw238):

8/3 = 2.67 8/9 = 0.89 to be more exact

OpenStudy (anonymous):

yes, unless the teacher is a jerk about it, i feel the easiest way is to keep it in fraction form

OpenStudy (cwrw238):

unless they asked for decimal form fractions are ok

OpenStudy (anonymous):

so how should i word my answer

OpenStudy (anonymous):

....ha, more politely put, yes

OpenStudy (anonymous):

show off and put decimal and fraction forms!!

OpenStudy (cwrw238):

yea why not!!

OpenStudy (anonymous):

hey @cwrw238, are you good with interests?

OpenStudy (anonymous):

so just divide each number by 3 for rule and 8, 2.67, 0.89

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this one was kinda simple. I worked it backwards. In my head, I asked myself what multiplies 24 to get 72. then 72 to get 216...